# TopoStreamer: Temporal Lane Segment Topology Reasoning in Autonomous Driving

### v2

@@ Introduction
where $\tilde{\textbf{Q}}$ denotes the updated queries by this layer. Then, the reference points are refined through offset adjustments to enable more precise sampling. This refinement facilitates the injection of more accurate positional embeddings into the queries in subsequent decoder layers, thereby enhancing the learning of precise lane segment localization. For more implementation details regarding the decoder process, please refer to the supplementary materials.
Figure 4: DN query generator for lane segment denoising.
Lane Segment Denoising
Lane topology reasoning, compared to map perception, requires detecting more fragmented lane segments. This is because shorter and more refined lane segments exhibit more complex topological connections, which are crucial for flexible and adaptive autonomous driving planning. However, accurately predicting topological relationships between these refined segments remains a significant challenge, while the temporal propagation of multiple lane segments’ attribute classification further compounds the complexity of the prediction task. Inspired by (Li et al. 2022a; Wang et al. 2024b), we design denoising objectives specifically for position estimation, topological consistency, and category attribute prediction of lane segments.
@@
Subsequently, we design some denoising loss to rectify contaminated coordinates, misclassified categories, and erroneous topological relationships caused by coordinate deviations in the DN queries. The denoising losses include L1 loss for position regression, cross-entropy and focal loss (Lin et al. 2017b) for classification and adjacency matrix prediction:
@@
where, for brevity, we omit the weights for each loss term. Details of DN losses can be found in supplementary materials. This module enhances the learning of diverse patterns in lane segments, specifically correcting erroneous predictions caused by temporal projection errors.
Training Loss
@@
where the lane segment loss $\mathcal{L}_{ls}$ supervises predicted lane segments through Hungarian matching (Li et al. 2023b), while $\mathcal{L}_{stream}$ and $\mathcal{L}_{denoise}$ are loss specifically optimized for lane segments streaming and denoising. $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ are hyperparameters.
Figure 5: The temporally accumulated perception results are compared across GT, SQD-MapNet (Wang et al. 2024b), and our TopoStreamer. Missing detections are highlighted with red color. For better viewing, zoom in on the image.
Experiments
We evaluate our method on multi-view lane topology benchmark OpenLane-V2 (Wang et al. 2024a). Since lane segment labels are exclusively available in subsetA, our validation is primarily conducted on this subset.
Method Epochs Temporal mAP $\uparrow$ $\text{AP}_{ls}\uparrow$ $\text{AP}_{ped}\uparrow$ $\text{TOP}_{lsls}\uparrow$ $\text{Acc}_{b}\uparrow$ FPS
MapTR (Liao et al. 2022) 24 No 27.0 25.9 28.1 - - 14.5
MapTRv2 (Liao et al. 2023) 24 No 28.5 26.6 30.4 - - 13.6
TopoNet (Li et al. 2023a) 24 No 23.0 23.9 22.0 - - 10.5
LaneSegNet (Li et al. 2023b) 24 No 32.6 32.3 32.9 25.4 45.9 14.7
TopoLogic (Fu et al. 2025a) 24 No 33.2 33.0 33.4 30.8 - -
Topo2Seq (Yang et al. 2025a) 24 No 33.6 33.7 33.5 26.9 48.1 14.7
StreamMapNet (Yuan et al. 2024) 24 Yes 20.3 22.1 18.6 13.2 33.2 14.1
SQD-MapNet (Wang et al. 2024b) 24 Yes 26.0 27.1 24.9 16.6 39.4 14.1
TopoStreamer (ours) 24 Yes 36.6 $\uparrow 3.0\%$ 35.0 $\uparrow 1.3\%$ 38.1 $\uparrow 4.6\%$ 28.5 $\downarrow 2.3\%$ 50.0 $\uparrow 1.9\%$ 13.6
Table 1: Comparison with the state-of-the-arts on OpenLane-V2 benchmark on lane segment. While our topology metrics do not surpass those of TopoLogic, our improvements in other metrics are substantial. Although we can integrate TopoLogic’s dedicated topology measures, our emphasis lies in planning-related comprehension. To ensure a fair comparison with other approaches, we do not include them.
@@
VectorMapNet (Liu et al. 2023) 24 No 13.8 11.1 2.7
STSU (Can et al. 2021) 24 No 14.9 12.7 2.9
MapTR (Liao et al. 2022) 24 No 21.0 17.7 5.9
TopoNet (Li et al. 2023a) 24 No 30.8 28.6 10.9
Topo2D (Li et al. 2024a) 24 No 38.2 29.1 26.2
TopoMLP (Wu et al. 2023) 24 No 37.4 28.3 21.7
LaneSegNet (Li et al. 2023b) 24 No 40.7 31.1 25.3
Topologic (Fu et al. 2025a) 24 No 39.4 29.9 23.9
Topo2Seq (Yang et al. 2025a) 24 No 42.7 33.5 27.0
TopoFormer (Lv et al. 2025) 24 No 42.1 34.7 24.7
StreamMapNet (Yuan et al. 2024) 24 Yes 28.8 21.7 12.9
SQD-MapNet (Wang et al. 2024b) 24 Yes 33.9 27.2 16.4
TopoStreamer (ours) 24 Yes 44.4 $\uparrow 1.7\%$ 35.2 $\uparrow 0.5\%$ 28.8 $\uparrow 1.8\%$
Table 2: Comparison with the state-of-the-arts on OpenLane-V2 benchmark on centerline perception. TopoFormer adopts a staged training strategy that utilizes a pretrained lane detector for topology reasoning training. While this leads to better detection performance, it offers only slight advantage in topology prediction.
Method mAP APls APped TOPlsls Accb
BL(w. Static PE) 32.6 32.3 32.9 25.4 45.9
+DBPE 33.5 $\uparrow 0.9\%$ 32.6 $\uparrow 0.3\%$ 34.3 $\uparrow 1.4\%$ 26.1 $\uparrow 0.7\%$ 47.2 $\uparrow 1.3\%$
+DBPE+DN 34.5 $\uparrow 1.9\%$ 32.9 $\uparrow 0.6\%$ 36.1 $\uparrow 3.2\%$ 26.7 $\uparrow 1.3\%$ 47.6 $\uparrow 1.7\%$
+Stream+Static PE 33.8 $\uparrow 1.2\%$ 32.8 $\uparrow 0.5\%$ 34.8 $\uparrow 1.9\%$ 26.3 $\uparrow 0.9\%$ 47.7 $\uparrow 1.8\%$
+Stream w/o PE 33.5 $\uparrow 0.9\%$ 32.0 $\downarrow 0.3\%$ 34.9 $\uparrow 2.0\%$ 27.1 $\uparrow 1.7\%$ 48.3 $\uparrow 2.4\%$
+Stream+DBPE 35.6 $\uparrow 3.0\%$ 34.7 $\uparrow 2.4\%$ 36.5 $\uparrow 3.6\%$ 27.8 $\uparrow 2.4\%$ 49.5 $\uparrow 3.6\%$
+Stream+DBPE+DN w/o IDTrack 34.4 $\uparrow 1.8\%$ 33.6 $\uparrow 1.3\%$ 36.2 $\uparrow 3.3\%$ 28.0 $\uparrow 2.6\%$ 49.5 $\uparrow 3.6\%$
+Stream+DBPE+DN 36.6 $\uparrow 4.0\%$ 35.0 $\uparrow 2.7\%$ 38.1 $\uparrow 5.2\%$ 28.5 $\uparrow 3.1\%$ 50.0 $\uparrow 4.1\%$
Table 3: Ablation study on different modules in our method. Options that enhance performance are highlighted in green, while those that degrade performance are marked in red.
@@
Table 4: Ablation studies on streaming attribute constraints.
Datasets and Metrics
OpenLane-V2 (Wang et al. 2024a) is a widely-used dataset for lane topology reasoning. Its subsetA, re-annotated from Argoverse 2 (Wilson et al. 2023), provides enhanced details on traffic signals, centerlines, lane boundaries, and their topological relationships. This subset includes over 20,000 training frames and more than 4,800 validation frames, with each frame comprising 7 camera images at a resolution of 2048 $\times$ 1550.
Metrics. We evaluate our model on two tasks: lane segment and centerline perception. The lane quality are evaluated under Chamfer distance and Frechet distance under a preset thresholds of {1.0, 2.0, 3.0} meters. For lane segment, mAP computed as average of APls and APped. APls and APped are used to estimate the quality of lane segment of road lines and pedestrian crossing, respectively. TOPlsls measures the performance of topology reasoning. We design a new metric Accb to evaluate lane boundary classification accuracy, which can be referred in supplementary materials. The metrics in cenerline perception are similar with those in lane segment. OLS (Wang et al. 2024a) is calculated between DETl and TOPll.
Experimental Settings
We adopt a pre-trained ResNet-50 (He et al. 2016), FPN (Lin et al. 2017a) and BevFormer (Li et al. 2022c) to encode the images to BEV features. The BEV grid is 200 $\times$ 100, which the perception range is $\pm$ 50m $\times$ $\pm$ 25m. Our decoder is based on Deformable DETR, with the cross-attention replaced by lane attention (Li et al. 2023b). The number of layer is 6. We use 200 queries, with 30% allocated for temporal propagation. The centerline, left boundary line, and right boundary line are each represented as individual sets of 10 ordered points in our predictions. We select 8 boundary reference points (4 from the left boundary and 4 from the right) to generate PE. The number of DN groups is dynamically adjusted based on batch instances, while DN queries are fixed at 240. Positional noise is introduced via box shifting (Wang et al. 2024b) with a factor of 0.2, and labels have a 50% flip probability. Training is conducted for 24 epochs with a batch size of 8 on NVIDIA A100 GPUs, with the first 12 epochs using single-frame input to stabilize streaming training. The initial learning rate is $2\times 10^{-4}$ with a cosine annealing schedule during training. AdamW (Kingma and Ba 2015) is adopted as optimizer. The values of $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ are set to 1.0, 0.3 and 1.0, respectively.
We re-train StreamMapNet (Wang et al. 2024b) and SQD-MapNet (Wang et al. 2024b), both of which predict 10 points for the left and right boundaries. The centerline is obtained by calculating the average positions of two boundaries.
Main Results
Model Analysis
Module Ablations. The first row in Tab. 3 presents our baseline (BL) model, LaneSegNet. LaneSegNet injects static PE into the queries. With the introduction of dynamic lane boundary PE (DBPE), the model exhibits a slight improvement. Further enhancement is achieved by incorporating lane segment denoising. These results demonstrate that the incorporation of dynamic lane boundary PE and lane segment denoising effectively improves the overall performance of the per-frame detection model. Comparing row 6 with row 2, when the baseline model is adapted to the streaming paradigm and supervised with streaming attribute constraints, significant improvements are observed (35.6% v.s. 33.5% in mAP, 27.8% v.s. 26.1% in TOPlsls, and 49.5% v.s. 47.2% in Accb). However, substituting DBPE with either initial static PE or no PE at all adversely impacts performance, particularly resulting in a 2% reduction in mAP. Finally, the addition of denoising leads to optimal performance. However, as demonstrated in the row 7, the model exhibits a significant decline in detection performance when the unique IDs of positive instances are not tracked within streaming attribute constraints to ensure lossless supervision.
Ablation Studies for Streaming Attribute Constraints. The results are shown in Tab. 4. The baseline implementation, corresponding to the first row, incorporates the DBPE module into the streaming framework while excluding both the streaming attribute constraints and lane segment denoising components. Introducing class constraint in streaming can achieve a considerable improvement in lane boundary classification. Subsequently, the progressive integration of mask and coordinate constraints leads to enhanced detection capability and improved topology reasoning performance.
Ablation Studies for Lane Segment Denoising. The results are presented in Tab. 5. The baseline implementation remains consistent with Tab. 4. Incorporating class denoising for content information in queries enhances performance in both detection and lane boundary classification, highlighting the importance of content learning in perception. Additionally, topology denoising enhances the robustness of topological reasoning against coordinate noise, while coordinate denoising significantly boosts detection performance.
Qualitative Results
As shown in Fig. 5, TopoStreamer can predict a complete road network with clear lane boundaries and accurate topology connections. More qualitative results can be found in the supplementary materials.
@@
Table 5: Ablation studies on lane segment denoising.
Conclusion

### v3

@@ 3.3 Dynamic Lane Boundary PE
where $\tilde{\textbf{Q}}$ denotes the updated queries by this layer. Then, the reference points are refined through offset adjustments to enable more precise sampling. This refinement facilitates the injection of more accurate positional embeddings into the queries in subsequent decoder layers, thereby enhancing the learning of precise lane segment localization. For more implementation details regarding the decoder process, please refer to the appendix.
3.4 Lane Segment Denoising
Lane topology reasoning, compared to map perception, requires detecting more fragmented lane segments. This is because shorter and more refined lane segments exhibit more complex topological connections, which are crucial for flexible and adaptive autonomous driving planning. However, accurately predicting topological relationships between these refined segments remains a significant challenge, while the temporal propagation of multiple lane segments’ attribute classification further compounds the complexity of the prediction task. Inspired by (Li et al., 2022a; Wang et al., 2024b), we design denoising objectives specifically for position estimation, topological consistency, and category attribute prediction of lane segments.
Figure 4: DN query for lane segment denoising.
@@ 3.4 Lane Segment Denoising
Subsequently, we introduce a set of denoising losses to rectify noisy coordinates, category misclassifications, and topological errors arising from coordinate deviations in the DN queries. The denoising losses include L1 loss for position regression, cross-entropy and focal loss (Lin et al., 2017b) for classification and adjacency matrix prediction:
@@
where, for brevity, we omit the weights for each loss term. Details of DN losses can be found in appendix. This module enhances the learning of diverse patterns in lane segments, specifically correcting erroneous predictions caused by temporal projection errors.
3.5 Training Loss
@@
where the lane segment loss $\mathcal{L}_{ls}$ supervises predicted lane segments through Hungarian matching (Li et al., 2023b), while $\mathcal{L}_{stream}$ and $\mathcal{L}_{denoise}$ are loss specifically optimized for lane segments streaming and denoising. $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ are hyperparameters.
4 Experiments
We evaluate our method on multi-view lane topology benchmark OpenLane-V2 (Wang et al., 2024a). Since lane segment labels are exclusively available in subsetA, our validation is primarily conducted on this subset. The results on subsetB can be found in appendix.
4.1 Datasets and Metrics
OpenLane-V2 (Wang et al., 2024a) is a widely-used dataset for lane topology reasoning. Its subsetA, re-annotated from Argoverse 2 (Wilson et al., 2023), provides enhanced details on traffic signals, centerlines, lane boundaries, and their topological relationships. This subset includes over 20,000 training frames and more than 4,800 validation frames, with each frame comprising 7 camera images at a resolution of 2048 $\times$ 1550.
Metrics. We evaluate our model on two tasks: lane segment and centerline perception. The lane quality are evaluated under Chamfer distance and Frechet distance under a preset thresholds of {1.0, 2.0, 3.0} meters. For lane segment, mAP computed as average of APls and APped. APls and APped are used to estimate the quality of lane segment of road lines and pedestrian crossing, respectively. TOPlsls measures the performance of topology reasoning. We design a new metric Accb to evaluate lane boundary classification accuracy, which can be referred in appendix. The metrics in cenerline perception are similar with those in lane segment. OLS (Wang et al., 2024a) is calculated between DETl and TOPll.
Table 1: Comparison with the state-of-the-arts on OpenLane-V2 benchmark on lane segment. All models adopt ResNet-50 as the backbone network and are trained for 24 epochs. Although we can integrate TopoLogic’s dedicated topology measures to enhance Toplsls, we do not include them for a fair comparison with other approaches.
Method Venue Temporal mAP $\uparrow$ $\text{AP}_{ls}\uparrow$ $\text{AP}_{ped}\uparrow$ $\text{TOP}_{lsls}\uparrow$ $\text{Acc}_{b}\uparrow$ FPS
MapTR (Liao et al., 2022) ICLR23 No 27.0 25.9 28.1 - - 14.5
MapTRv2 (Liao et al., 2023) IJCV24 No 28.5 26.6 30.4 - - 13.6
TopoNet (Li et al., 2023a) Arxiv23 No 23.0 23.9 22.0 - - 10.5
LaneSegNet (Li et al., 2023b) ICLR24 No 32.6 32.3 32.9 25.4 45.9 14.7
TopoLogic (Fu et al., 2025a) NIPS24 No 33.2 33.0 33.4 30.8 - -
Topo2Seq (Yang et al., 2025a) AAAI25 No 33.6 33.7 33.5 26.9 48.1 14.7
StreamMapNet (Yuan et al., 2024) WACV24 Yes 20.3 22.1 18.6 13.2 33.2 14.1
SQD-MapNet (Wang et al., 2024b) ECCV24 Yes 26.0 27.1 24.9 16.6 39.4 14.1
@@ 4.1 Datasets and Metrics
Figure 5: Comparison with the state-of-the-arts on OpenLane-V2 benchmark on centerline perception. All models adopt ResNet-50 as the backbone network and are trained for 24 epochs. Unlike other methods, TopoFormer adopts a staged training strategy that utilizes a pretrained lane detector for topology reasoning training. While this leads to better detection performance, it offers only slight advantage in topology prediction.
Method Venue Temporal $\text{OLS}\uparrow$ $\text{DET}_{l}\uparrow$ $\text{TOP}_{ll}\uparrow$
VectorMapNet (Liu et al., 2023) ICML23 No 13.8 11.1 2.7
STSU (Can et al., 2021) ICCV21 No 14.9 12.7 2.9
MapTR (Liao et al., 2022) ICLR23 No 21.0 17.7 5.9
TopoNet (Li et al., 2023a) Arxiv23 No 30.8 28.6 10.9
Topo2D (Li et al., 2024a) Arxiv24 No 38.2 29.1 26.2
TopoMLP (Wu et al., 2023) ICLR24 No 37.4 28.3 21.7
LaneSegNet (Li et al., 2023b) ICLR24 No 40.7 31.1 25.3
TopoLogic (Fu et al., 2025a) NIPS24 No 39.4 29.9 23.9
Topo2Seq (Yang et al., 2025a) AAAI25 No 42.7 33.5 27.0
TopoFormer (Lv et al., 2025) CVPR25 No 42.1 34.7 24.7
StreamMapNet (Yuan et al., 2024) WACV24 Yes 28.8 21.7 12.9
SQD-MapNet (Wang et al., 2024b) ECCV24 Yes 33.9 27.2 16.4
[…]
4.2 Experimental Settings
We adopt a pre-trained ResNet-50 (He et al., 2016), FPN (Lin et al., 2017a) and BevFormer (Li et al., 2022c) to encode the images to BEV features. The BEV grid is 200 $\times$ 100, which the perception range is $\pm$ 50m $\times$ $\pm$ 25m. Our decoder is based on Deformable DETR, with the cross-attention replaced by lane attention (Li et al., 2023b). The number of layer is 6. We use 200 queries, with 30% allocated for temporal propagation. The centerline, left boundary line, and right boundary line are each represented as individual sets of 10 ordered points in our predictions. We select 8 boundary reference points (4 from the left boundary and 4 from the right) to generate PE. The number of DN groups is dynamically adjusted based on batch instances, while DN queries are fixed at 240. Positional noise is introduced via box shifting (Wang et al., 2024b) with a factor of 0.2, and labels have a 50% flip probability. Training is conducted for 24 epochs with a batch size of 8 on NVIDIA V100 GPUs, with the first 12 epochs using single-frame input to stabilize streaming training. The initial learning rate is $2\times 10^{-4}$ with a cosine annealing schedule during training. AdamW (Kingma & Ba, 2015) is adopted as optimizer. The values of $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ are set to 1.0, 0.3 and 1.0, respectively. The confidence threshold for the adjacency matrix is set at 0.5, and all visualized segments must exceed a threshold of 0.3.
We re-train StreamMapNet (Wang et al., 2024b) and SQD-MapNet (Wang et al., 2024b), both of which predict 10 points for the left and right boundaries. The centerline is obtained by calculating the average positions of two boundaries.
4.3 Main Results
Figure 6: The temporally accumulated perception results are compared across GT, our TopoStreamer, and SQD-MapNet (Wang et al., 2024b). The images from top to bottom correspond to GT, TopoStreamer, and SQD-MapNet. Missing detections are highlighted with red color. For better viewing, zoom in on the image.
Table 2: Ablation study results on losses.
@@ 4.3 Main Results
(a) Ablation studies on streaming attribute constraints.
@@
(b) Ablation studies on lane segment denoising.
4.4 Model Analysis
Ablation Studies for Streaming Attribute Constraints. The results are shown in Tab. 2(a). The baseline implementation, corresponding to the first row, incorporates the DBPE module into the streaming framework while excluding both the streaming attribute constraints and lane segment denoising components. Introducing class constraint in streaming can achieve a considerable improvement in lane boundary classification. Subsequently, the progressive integration of mask and coordinate constraints leads to enhanced detection capability and improved topology reasoning performance.
Ablation Studies for Lane Segment Denoising. The results are presented in Tab. 2(b). The baseline implementation remains consistent with Tab. 2(a). Incorporating class denoising for content information in queries enhances performance in both detection and lane boundary classification, highlighting the importance of content learning in perception. Additionally, topology denoising enhances the robustness of topological reasoning against coordinate noise, while coordinate denoising boosts detection performance.
Figure 7: Ablation study on different modules in our method.
Method mAP APls APped TOPlsls Accb
BL(w. Static PE) 32.6 32.3 32.9 25.4 45.9
+DBPE 33.5 32.6 34.3 26.1 47.2
+DBPE+DN 34.5 32.9 36.1 26.7 47.6
+Stream+Static PE 33.8 32.8 34.8 26.3 47.7
+Stream w/o PE 33.5 32.0 34.9 27.1 48.3
+Stream+DBPE 35.6 34.7 36.5 27.8 49.5
+Stream+DBPE+DN w/o IDTrack 34.4 33.6 36.2 28.0 49.5
+Stream+DBPE+DN 36.6 35.0 38.1 28.5 50.0
Module Ablations. The first row in Tab. 7 presents our baseline (BL) model, LaneSegNet. LaneSegNet injects static PE into the queries. With the introduction of dynamic lane boundary PE (DBPE), the model exhibits a slight improvement. Further enhancement is achieved by incorporating lane segment denoising. These results demonstrate that the incorporation of dynamic lane boundary PE and lane segment denoising effectively improves the overall performance of the per-frame detection model. Comparing row 6 with row 2, when the baseline model is adapted to the streaming paradigm and supervised with streaming attribute constraints, considerable improvements are observed (35.6% v.s. 33.5% in mAP, 27.8% v.s. 26.1% in TOPlsls, and 49.5% v.s. 47.2% in Accb). However, substituting DBPE with either initial static PE or no PE at all adversely impacts performance, particularly resulting in a 2% reduction in mAP. Finally, the addition of denoising leads to optimal performance. However, as demonstrated in the row 7, the model exhibits a significant decline in detection performance when the unique IDs of positive instances are not tracked within streaming attribute constraints to ensure lossless supervision.
Figure 8: Qualitative results under different road structures. The images from top to bottom correspond to GT, our TopoStreamer, and LaneSegNet(Li et al., 2023b). For better viewing, zoom in on the image.
(a) Results when the ego vehicle is moving.
(b) Results when the ego vehicle is stationary.
Figure 9: Visualization of topology predictions across consecutive 5 frames. The results of TopStreamer are shown on the top, and the results of LaneSegNet (Li et al., 2023b) are shown on the bottom. For better viewing, zoom in on the images. This demonstrates TopoStreamer’s capability for accurate detection with temporal consistency.
4.5 Qualitative Results
As shown in Fig. 6, 8 and 9, TopoStreamer is capable of predicting a complete road network with clearly lane boundaries, accurate topology connections, and temporal consistency. Additional qualitative results are provided in the appendix.
5 Conclusion
@@
Appendix A Appendix
A.1 Streaming Attribute Constraints.
We employ MLPs to predict lane segment coordinate, lane segment class, boundary class and BEV mask from stream queries $Q_{t}^{S}$:
$\begin{split}\tilde{\mathbf{L}}_{t}^{c}&=\text{MLP}_{reg}(Q_{t}^{S})+\text{InSigmod}(\mathbf{R^{S}_{C}})\\ \tilde{\mathbf{L}}_{t}^{c}&=\text{Denorm}(\text{sigmoid}(\tilde{\mathbf{L}}_{t}^{c}))\\ offset&=\text{MLP}_{offset}(Q_{t}^{S})\\ \tilde{\mathbf{L}}_{t}^{l}&=\tilde{\mathit{L}}_{t}^{c}+offset,\tilde{\mathbf{L}}_{t}^{r}=\tilde{\mathbf{L}}_{t}^{c}-offset\\ \tilde{\mathbf{L}}_{t}&=\text{Concat}(\tilde{\mathbf{L}}_{t}^{c},\tilde{\mathbf{L}}_{t}^{l},\tilde{\mathbf{L}}_{t}^{r})\\ \tilde{\mathit{C}}_{t}&=\text{MLP}_{cls}(Q_{t}^{S})\\ \tilde{\mathit{T}}_{t}&=\text{MLP}_{bcls}(Q_{t}^{S})\\ \tilde{\mathbf{M}}_{t}&=\text{Sigmoid}(\text{MLP}_{mask}(Q_{t}^{S})\otimes\tilde{\textbf{F}}_{bev}^{t})\end{split}$ (1)
where InSigmod refers to the inverse sigmoid function, while Denorm stands for denormalize. Then, the streaming attribute constraints are represented as:
$\begin{split}\mathcal{L}_{coord}^{Stream}&=\mathcal{L}_{L1}(\tilde{\mathbf{L}}_{t},\mathbf{L}_{t})\\ \mathcal{L}_{cls}^{Stream}&=\kappa_{1}\mathcal{L}_{Focal}(\tilde{\mathit{C}}_{t},\mathit{C}_{t})+\kappa_{2}\mathcal{L}_{CE}(\tilde{\mathit{T}}_{t},\mathit{T}_{t})\\ \mathcal{L}_{mask}^{Stream}&=\kappa_{3}\mathcal{L}_{CE}(\tilde{\mathbf{M}}_{t},\mathbf{M}_{t})+\kappa_{4}\mathcal{L}_{Dice}(\tilde{\mathbf{M}}_{t},\mathbf{M}_{t})\\ \mathcal{L}_{Stream}&=\kappa_{5}\mathcal{L}_{coord}^{Stream}+\kappa_{6}\mathcal{L}_{cls}^{Stream}+\kappa_{7}\mathcal{L}_{mask}^{Stream}\end{split}$ (2)
where the values of $\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$, $\kappa_{4}$, $\kappa_{5}$, $\kappa_{6}$, and $\kappa_{7}$ are 1.0, 0.01, 1.0, 1.0, 0.025, 1.0 and 3.0.
A.2 Lane Segment Denoising
Lane segment denoising applies controlled noise to annotations and then removes it, thereby improving the model’s capability to learn the diverse patterns present in lane segments. In the lane segment denoising, we predict position, classification and adjacency matrix from denoising (DN) queries $Q^{D}$ as follows:
[…]
Then, the denoising loss function is defined as:
$\begin{split}\mathcal{L}_{coord}^{DN}&=\mathcal{L}_{L1}(\tilde{\mathbf{L}},\mathbf{L})\\ \mathcal{L}_{cls}^{DN}&=\beta_{1}\mathcal{L}_{Focal}(\tilde{\mathit{C}},\mathit{C})+\beta_{2}\mathcal{L}_{CE}(\tilde{\mathit{T}},\mathit{T})\\ \mathcal{L}_{Topo}^{DN}&=\mathcal{L}_{Focal}(\tilde{\mathbf{A}},\mathbf{A})\\ \mathcal{L}_{denoise}&=\lambda_{1}\mathcal{L}_{coord}^{DN}+\lambda_{2}\mathcal{L}_{cls}^{DN}+\lambda_{3}\mathcal{L}_{topo}^{DN}\end{split}$ (4)
where the hyperparameters are defined as: $\beta_{1}=1.0$, $\beta_{2}=0.01$, $\lambda_{1}=0.025$, $\lambda_{2}=1.0$ and $\lambda_{3}=5.0$. Some examples of lane segment denoising are shown in Fig. 2. For better visualization, we only display the denoising results of the centerlines. It can be observed that the added noise disrupts the connectivity of the road network. Through the denoising process, the original positions and connectivity relationships are effectively restored. This enhances the model’s ability to predict both the positional and connectivity topology of lane segments.
A.3 Total Loss Function
The overall loss function in TopoSteamer is defined as:
where $\alpha_{1}=1.0$, $\alpha_{1}=0.3$ and $\alpha_{1}=1.0$, respectively. The lane segment loss is defined as:
$\mathcal{L}_{ls}=\omega_{1}\mathcal{L}_{vec}+\omega_{2}\mathcal{L}_{seg}+\omega_{3}\mathcal{L}_{cls}+\omega_{4}\mathcal{L}_{type}+\omega_{5}\mathcal{L}_{topo}$ (6)
where $\mathcal{L}_{seg}=\omega_{6}\mathcal{L}_{ce}+\omega_{7}\mathcal{L}_{dice}$ consists of a Cross-Entropy loss and a Dice loss used to supervise the BEV semantic mask., and the hyperparameters are defined as: $\omega_{1}=0.025$, $\omega_{2}=3.0$, $\omega_{3}=1.5$, $\omega_{4}=0.01$, $\omega_{5}=5.0$, $\omega_{6}=1.0$ and $\omega_{7}=1.0$. $\mathcal{L}_{vec}$ is the L1 loss computed between the predicted vectorized lanes and the ground truth lanes. The classification losses $\mathcal{L}_{cls}$ and $\mathcal{L}_{type}$ are used for lane segment classification. $\mathcal{L}_{topo}$ is a focal loss applied to supervise the topological relationship prediction.
A.4 Streaming Training
We adopt the streaming training strategy for temporal fusion. For each training sequence, we randomly divide it into 2 splits at the start of each training epoch to foster more diverse data sequences. During inference, we use the entire sequences. Suppose a batch contains N samples, each from a different scene, read in chronological order. Temporal fusion is performed by determining whether the current data and the previously read data belong to the same scene. To facilitate temporal fusion, we introduce several memory modules, including stream query memory, stream BEV memory, and stream reference point memory, which store the predictions from the preceding frame.
A.5 Metric for Lane Boundary Classification
Previous approach (Li et al., 2023b) classify lane boundaries as dashed, solid, or non-visible, but they don’t measure how accurate these predictions are. This accuracy is crucial for self-driving cars when making lane-change decisions. To solve this, we introduce a new metric to to evaluate lane boundary classification accuracy. We follow the design of Toplsls metric (Wang et al., 2024a). We first build a projection between predictions and ground truth to preserve true positive instances, according to Fréchet distance. Then, we evaluate the accuracy of the left and right boundary types by comparing the predicted types with the GT types.
Figure 1: The detailed architecture of decoder.
A.6 Decoder Architecture
The detailed architecture of the decoder is shown in Fig. 1. In the first layer of the decoder, we utilize identical initialization (Li et al., 2023b) to produce the initial centerline reference points and boundary reference points from the initial position embedding. The initial queries, combined with DN queries, are fed into the self-attention (SA) and then augmented with the position embedding generated from the initial and DN boundary reference points. The sampling of boundary reference points is based on the heads-to-regions sampling method (Li et al., 2023b). Specifically, sampling is performed at symmetric offset positions on both sides of the centerline reference points. The position embedding is obtained by applying sinusoidal encoding to the coordinates of the reference point. These queries interact with BEV features through lane attention (Li et al., 2023b), employing a heads-to-regions sampling mechanism guided by the boundary reference points. At the outset of layer 1, we predict an updated offset to refine the initial and DN center reference points, along with a boundary offset to determine the boundary reference points by applying it to the center reference points. Meanwhile, in this layer, stream query, stream center reference points and stream boundary reference points are employed to replace the lowest confidence N-k queries and their reference points. Here, N represents the predefined total number of queries, set at 200, while K denotes the number of stream queries, which is 66, accounting for 30% of the total queries. The same updating procedure as in layer 0 is then applied to these queries. The updating process remains consistent and regular across layers 2 to 5.
Method Venue Epochs Temporal $\text{OLS}\uparrow$ $\text{DET}_{l}\uparrow$ $\text{TOP}_{ll}\uparrow$
VectorMapNet (Liu et al., 2023) ICML23 24 No - 3.5 -
STSU (Can et al., 2021) ICCV21 24 No - 8.2 -
MapTR (Liao et al., 2022) ICLR23 24 No - 15.2 -
TopoNet (Li et al., 2023a) Arxiv23 24 No 25.1 24.3 6.7
TopoMLP (Wu et al., 2023) ICLR24 24 No 36.2 26.6 19.8
LaneSegNet (Li et al., 2023b) ICLR24 24 No 38.7 27.5 24.9
TopoLogic (Fu et al., 2025a) NIPS24 24 No 36.2 25.9 21.6
StreamMapNet (Yuan et al., 2024) WACV24 24 Yes 26.7 18.9 11.9
SQD-MapNet (Wang et al., 2024b) ECCV24 24 Yes 29.1 21.9 13.3
[…]
Table 1: Comparison with the state-of-the-arts on OpenLane-V2 subsetB on centerline perception.
Figure 2: Qualitative results for lane segment denoising.
Top-K mAP Toplsls
10% 34.8 27.9
30% 36.6 28.5
50% 34.3 27.3
75% 32.9 25.1
(a) Different numbers of stream queries.
Number mAP Toplsls
120 35.0 27.4
240 36.6 28.5
360 35.1 27.7
(b) Different numbers of DN queries.
Table 2: Ablation study results.
A.7 Experiment
We provide comparative experiments on the OpenLane-V2 subset-B benchmark. In fact, this benchmark do not contain lane segment annotations with only centerline annotations. We generate pseudo-labels by augmenting the lane centerlines with a standardized lane width of 4 meters. The results are shown in Tab. 1. We outperforms LaneSegNet by 3.9% OLS, 3.4% DETl, and 4.5 % TOPll.
We present additional experiments focusing on the selection of the number of stream queries and DN queries on subset-A. The results of the ablation study investigating the impact of varying numbers of stream queries are presented in Tab. 2(a). Optimal performance is attained when 30% of the queries from the preceding frame are selected for temporal propagation. The results of the ablation study about the number of DN queries are shown in Tab. 2(b). Setting the number of DN queries to 240 yields the optimal performance.
A.8 DEMO
See the supplementary material vis.gif file for details.
A.9 Limitation and future work
Current lane topology reasoning methods are affected by the long-tail problem, exhibiting limited detection confidence in regions with excessive curvature or indistinct lane markings. Therefore, we plan to explore the use of Vision-Language Models (VLMs) to address this issue. By leveraging VLMs to interpret road structures and generate Chains of Thought (CoT), we aim to provide prior knowledge for lane topology reasoning and enhance interpretability. Additionally, we will investigate integrating road topology with end-to-end autonomous driving systems, using lane topology to constrain vehicle trajectory planning and improve safety.
A.10 Use of LLM
In this paper, large language model is used only for writing enhancement purposes.